OpenStudy (anonymous):

What is the value of b^2 - 4ac for the following equation? 2x^2 + 3x = -1

OpenStudy (anonymous):

Choices; 0 1 17

OpenStudy (anonymous):

???

OpenStudy (anonymous):

ax^2+bx+c

OpenStudy (anonymous):

ax^2+bx+c=0

OpenStudy (anonymous):

2x2 + 3x +1=ax^2+bx+c a=2 b=3 c=1

OpenStudy (anonymous):

So then what's the value for b^2 - 4ac ?

OpenStudy (anonymous):

Apparently the right answer is 1. But what's the solution to that?

OpenStudy (anonymous):

@cinar

OpenStudy (anonymous):

its 1. b is 3 a is 2 and c is 1. so \[3^2-4(2)(1)\]

OpenStudy (anonymous):

then you get 9-8 which is 1

OpenStudy (anonymous):

Ah. Makes sense. Thank you!!! :)